Solve given differential equations by separation of variables

dy/dx=y^3/x^2
1
Expert's answer
2020-08-13T17:33:18-0400

dy/dx =y3/x2

dy/y3=dx/x2

integrating

-2*y-2=-x-1+c1, c1 is an integration constant,

y-2=(x-1+c1)/2

y-2=x-1/2 +c2, c2 is an integration constant, c2=c1/2.

y=(x-1/2+c2)-1/2